# Dual process in the two-parameter Poisson–Dirichlet diffusion

### v5

@@ 3 The dual process
In this section we fully characterize a dual process for the two-parameter Poisson–Dirichlet diffusion induced by symmetric monomial functions. This is shown to be a pure-death continuous-time Markov chain on integer partitions which coincides with the process counting the number of surviving lineages of each type in a Kingman’s coalescent tree, going backward in time, conditional on a starting unlabelled partition of “leaves”. The dual process does not depend on $\alpha$. In particular, the block-counting process is the same as the known block-counting process dual to any reversible, neutral Wright–Fisher-type diffusion. The dual will then be used in Section 5 for deriving the transition density of the two-parameter diffusion, together with some additional results established in Section 4 on the sampling distribution $M_{n}(\eta)$ conditional on a partially observed partition.
The duality identity (3), under reasonably general conditions, can be verified through a similar identity involving the corresponding infinitesimal generators acting on appropriate test functions, i.e., $Lh(\cdot,d)(x)=Ah(x,\cdot)(d)$ where $L$ is the generator of $X_{t}$ and $A$ that of the dual. See [49], Proposition 1.2. To this end, note now that the family of functions $\{{P}_{\eta}(x),\eta\in\Gamma,{\min_{i}\eta_{i}>1}\}$, with $P_{\eta}$ as in (4), are a linear basis for the algebra which gives the core of $L_{\alpha,\theta}$. Cf. [61], Section 2. Here $\eta$ is unranked and these functions are intended as the their continuous extension from $\nabla_{\infty}$ to $\overline{\nabla}_{\infty}$. Hence ${P}_{(1)}(x)=1$, which for example implies that $1={P}_{(1)}(x){P}_{(1)}(x)={P}_{(2)}(x)+{P}_{(1,1)}(x)$, from which ${P}_{(1,1)}(x):=1-{P}_{(2)}(x)$. More generally, we would like to compute $L_{\alpha,\theta}$ on all $P_{\eta}$. This is allowed by the following lemma. Denote by $e_{i}$ the canonical vector in the $i$-th direction, and let $\eta-e_{i}$ when $\eta_{i}=1$ indicate $(\eta_{1},\ldots,\eta_{i-1},\eta_{i+1},\ldots,\eta_{d})$.
@@
Note in the above Lemma that since ${P}_{\eta}$ is symmetric in $\eta$, ${P}_{\eta-e_{i}}$ and ${P}_{\eta-e_{i}+e_{j}}$ are well defined. The next Lemma, key in identifying the dual process, makes use of (9) to show how the operator $L_{\alpha,\theta}$ acts on all functions ${P}_{\eta}$.
@@
Let $\eta$ be unranked. $L_{\alpha,\theta}$ is well defined on all ${P}_{\eta}$ in the system of equations (9), recursive on $a_{1}(\eta)$. In particular $L_{\alpha,\theta}1=0$, and if $n=|\eta|$ and $d=l(\eta)$,
@@
This was first proved by [61], Proposition 3.1. In the appendix we provide an independent proof based on (9). ∎
@@
Let $X$ be the diffusion process corresponding to $L_{\alpha,\theta}$, and let $\{\mathcal{D}_{t}\}_{t\geq 0}$ be a continuous-time death process on $\Gamma$, with rates $\lambda_{|\eta|}p^{\downarrow}(\eta,\omega)$, where $p^{\downarrow}(\eta,\omega):=\eta_{i}{a}_{\eta_{i}}(\eta)/|\eta|$ when $\omega$ is the descending arrangement of $\eta-e_{i}$ and zero elsewhere. Then, for every $\eta\in\Gamma$ and $x\in\overline{\nabla}_{\infty},$
(13) $\mathbb{E}\big{[}g_{\mathcal{D}_{t}}(x)|\mathcal{D}_{0}=\eta]=\mathbb{E}\big{[}g_{\eta}(X(t))|X_{0}=x].$
@@
Since $g_{\eta}(x)$ is symmetric with respect to $\eta$, from Lemma 3.2 we have
@@
If we now define $A_{\theta}$ to be the operator on the right hand side of (14) acting on $g_{\eta}(x)$ as a function of $\eta$, it is plain that $A_{\theta}$ defines a pure-death process $\mathcal{D}_{t}$ with rates as in the claim. The identity (13) now follows from Proposition 1.2 in [49] and the Feller property of both $X_{t}$ and $\mathcal{D}_{t}$. ∎
Here $\mathcal{D}_{t}$ is the process describing the group sizes in Kingman’s coalescent with mutation at time $t$. The next proposition identifies its transition probability function. To this end, let
@@
${\cal H}(\omega|\eta)$ is the probability of obtaining an unlabelled colour configuration $\omega$ when sampling $|\omega|$ balls without replacement uniformly from an urn containing a colour configuration of $\eta$ balls. The $|\omega|$ balls may be sampled all at once or one by one.
@@
(19) $q^{\theta}_{\eta\omega}(t):=\mathbb{P}(\mathcal{D}_{t}=\omega|\mathcal{D}_{0}=\eta)={\cal H}(\omega|\eta)d^{\theta}_{|\eta||\omega|}(t),\quad\quad\omega\subset\eta,$
and zero otherwise, with $d^{\theta}_{nm}(t)$ for $n\geq m\geq 1$ as in (16), and with $1$ as absorbing state.
@@
$d_{\eta\omega}^{\theta}(t)=$ $\,\mathbb{P}_{\eta}(\mathcal{D}_{t}=\omega)=\mathbb{P}_{\eta}(\mathcal{D}_{t}=\omega|D_{t}=|\omega|)\mathbb{P}_{\eta}(D_{t}=|\omega|)$
@@ 4 Two-parameter conditional partition structures
Let $w=|\omega|$, and suppose an urn contains a single white ball and $w$ non-white balls, whose colours are denoted by $Y^{w}=(Y_{1},\cdots,Y_{w})$, with $r\leq w$ different colours. Define $\pi:Y^{w}\to\omega$ to be the function that maps the sample $Y^{w}$ into the partition $\omega=(\omega_{1},\cdots,\omega_{r})\in\Gamma_{w}$, determined by the equivalence relation on colours, whereby $Y_{i},Y_{j}$ are in the same group if $Y_{i}=Y_{j}$. For every group $\omega_{j}$, one ball is assigned mass $1-\alpha$ and the remaining $\omega_{j}-1$ balls are assigned mass $1$. The single white ball in the urn is assigned mass $\theta+\alpha r$. Balls are then drawn sequentially as follows. If the white ball is drawn, a ball of a new colour is added to the urn and assigned mass $1-\alpha$, and the mass of the white ball is increased by $\alpha$. If a coloured ball is drawn, it is replaced in the urn together with an additional ball of mass $1$ of the same colour. This urn model is an extension of (2) and [46].
When the total number of balls in the urn is $n$, denote the colours of the $n-w$ new balls by $X^{n-w}=(X_{1},\ldots,X_{n-w})$. Let the colour configuration be denoted by $\pi(X^{n-w})$ for the new sample and $\pi(X^{n-w},Y^{w})$ for the combined sample. We are interested in the distribution of these colour configurations conditional on $Y^{w}$. To this end, denote by $\text{Dir}(\beta_{1},\ldots,\beta_{r})$ a Dirichlet distribution with parameters $(\beta_{1},\ldots,\beta_{r})$ on the $(r-1)$-dimensional simplex, and let a direct sum notation $\oplus$ indicate an accumulation of the points of two point processes. Finally, denote by $\mathrm{PD}_{\alpha,\theta|\omega}$ a $\mathrm{PD}_{\alpha,\theta}$ distribution conditional on having observed the partition $\omega$, and by $\mathbb{E}_{\alpha,\theta|\omega}$ the corresponding expectation.
@@
Let $\omega\in\Gamma_{w},\eta\in\Gamma_{n}$ and $\gamma\in\Gamma_{n-w}$. Then $\mathrm{PD}_{\alpha,\theta|\omega}$ equals in distribution
(21) $Z_{w,r}\mathrm{Dir}(\omega_{1}-\alpha,\ldots,\omega_{r}-\alpha)\oplus(1-Z_{w,r})\mathrm{PD}_{\alpha,\theta+\alpha r},$
where $Z_{w,r}$ has Beta distribution $\mathrm{Beta}(w-r\alpha,\theta+r\alpha)$, independent of everything else. Furthermore,
(22) $\mathbb{P}(\pi(X^{n-w})=\gamma|\pi(Y^{w})=\omega)=$ $\,\mathbb{E}_{\alpha,\theta|\omega}\big{[}\vec{P}_{\gamma}\big{]},$
$\mathbb{P}(\pi(X^{n-w},Y^{w})=\eta|\pi(Y^{w})=\omega)=$ $\,{\cal H}(\omega|\eta)\frac{\mathbb{E}_{\alpha,\theta}\big{[}\vec{P}_{\eta}\big{]}}{\mathbb{E}_{\alpha,\theta}\big{[}\vec{P}_{\omega}\big{]}},$
@@
Equations (21) and (22) follow from the exchangeability of draws in the above described urn model, together with Corollary 20 in [63]. Consider now a path from $\omega$ to $\eta$. Let $\omega^{i}=\pi(X^{i},Y^{w}),1\leq i\leq n-w$, $w^{i}=|\omega^{i}|$ and $r^{i}$ be the number of colours in $\omega^{i}$. Set $\omega^{0}=\omega$. When $\pi(X^{n-w},Y^{w})=\eta$, there is a path $\omega=\omega^{0}\subset\cdots\subset\omega^{n-w}=\eta$. Let $\chi_{B}(\omega^{i},\omega^{i+1})=a_{\omega^{i}_{k}}(\omega^{i})$ if $\omega^{i+1}$ is obtained from $\omega^{i}$ by adding $1$ to an existing $k$-th component or $\chi_{B}(\omega^{i},\omega^{i+1})=1$ if a new component is added to $\omega^{i+1}$. Then
$\mathbb{P}(\omega^{i+1}|\omega^{i})=\chi_{B}(\omega^{i},\omega^{i+1})\frac{\mathbb{E}_{\alpha,\theta}\big{[}{P}_{\omega^{i+1}}\big{]}}{\mathbb{E}_{\alpha,\theta}\big{[}{P}_{\omega^{i}}\big{]}}=\begin{cases}\displaystyle{\small a_{\omega_{k}^{i}}(\omega^{i})}\frac{\omega_{k}^{i}-\alpha}{\theta+w^{i}},\quad&\text{if a coloured ball is drawn,}\\[8.53581pt] \displaystyle\frac{\theta+\alpha r^{i}}{\theta+w^{i}},&\text{if the white ballis drawn.}\end{cases}$
@@
$\mathbb{P}(\eta|\omega)=$ $\,\sum_{\omega=\omega^{0}\subset\cdots\subset\omega^{n-w}=\eta}\prod_{i=0}^{n-w-1}\chi_{B}(\omega^{i},\omega^{i+1})\frac{\mathbb{E}_{\alpha,\theta}\big{[}{P}_{\omega^{i+1}}\big{]}}{\mathbb{E}_{\alpha,\theta}\big{[}{P}_{\omega^{i}}\big{]}}$
$=$ $\,\left(\sum_{\omega=\omega^{0}\subset\cdots\subset\omega^{n-w}=\eta}\prod_{i=0}^{n-w-1}\chi_{B}(\omega^{i},\omega^{i+1})\right)\frac{\mathbb{E}_{\alpha,\theta}\big{[}{P}_{\eta}\big{]}}{\mathbb{E}_{\alpha,\theta}\big{[}{P}_{\omega}\big{]}}$
$=$ $\,\frac{a_{1}(\omega)!\cdots a_{w}(\omega)!\dim(\omega,\eta)}{a_{1}(\eta)!\cdots a_{n}(\eta)!}\frac{\mathbb{E}_{\alpha,\theta}\big{[}{P}_{\eta}\big{]}}{\mathbb{E}_{\alpha,\theta}\big{[}{P}_{\omega}\big{]}}$
$=$ $\,\frac{\binom{w}{\omega}\dim(\omega,\eta)}{\binom{n}{\eta}}\frac{\binom{n}{\eta}\frac{1}{a_{1}(\eta)!\cdots a_{n}(\eta)!}\mathbb{E}_{\alpha,\theta}\big{[}{P}_{\eta}\big{]}}{\binom{w}{\omega}\frac{1}{a_{1}(\omega)!\cdots a_{w}(\omega)!}\mathbb{E}_{\alpha,\theta}\big{[}{P}_{\omega}\big{]}}={\cal H}(\omega|\eta)\frac{\mathbb{E}_{\alpha,\theta}\big{[}\vec{P}_{\eta}\big{]}}{\mathbb{E}_{\alpha,\theta}\big{[}\vec{P}_{\omega}\big{]}}$
@@
Note now that when $\gamma$ gives the counts in the new sample and $\eta$ the counts in the combined sample, we have $\mathbb{P}(\pi(X^{n-w})=\gamma|Y^{w})=\mathbb{P}(\pi(X^{n-w},Y^{w})=\eta|Y^{w})$ and therefore
@@
Let $V$ have distribution $\mathrm{PD}_{\alpha,\theta}$ and $Y^{w}$ be a sample from the urn. By Lemma 2.1, $\mathbb{P}(\pi(Y^{w})=\omega,V\in dy)=\vec{P}_{\omega}(y)\mathrm{PD}_{\alpha,\theta}(dy).$ Since the marginal distribution are $\mathbb{P}(\pi(Y^{w})=\omega)=\mathbb{E}_{\alpha,\theta}\big{[}\vec{P}_{\omega}\big{]}$, by Bayes’ theorem
$\mathbb{P}(V\in dy|\pi(Y^{w})=\omega)=\frac{\vec{P}_{\omega}(y)}{\mathbb{E}_{\alpha,\theta}\big{[}\vec{P}_{\omega}\big{]}}\mathrm{PD}_{\alpha,\theta}(dy).$
From (22) in Proposition 4.1, $\mathbb{P}(V\in dy|\pi(Y^{w})=\omega)=\mathrm{PD}_{\alpha,\theta|\omega}(dy)$, leading to the result. ∎

### v6

@@ 3 The dual process
In this section we fully characterize a dual process for the two-parameter diffusion induced by symmetric monomial functions. This is shown to be a pure-death continuous-time Markov chain on integer partitions which coincides with the process counting the number of surviving lineages of each type in a Kingman’s coalescent tree, going backward in time, conditional on a starting unlabelled partition of “leaves”. The dual process does not depend on $\alpha$ through its transition rates, but only through the duality functions. In particular, the block-counting process is the same as the known block-counting process dual to any reversible, neutral Wright–Fisher-type diffusion. The dual we describe in this Section will be used in Section 5 for deriving the transition density of the two-parameter diffusion, after providing some additional results in Section 4 on the sampling distribution $M_{n}(\eta)$ conditional on a partially observed partition.
The duality identity (3), under reasonably general conditions, can be verified through a similar identity involving the corresponding infinitesimal generators acting on appropriate test functions, which is the object of the next two Lemmas, i.e., $Lh(\cdot,d)(x)=Ah(x,\cdot)(d)$ where $L$ is the generator of $X_{t}$ and $A$ that of the dual. See [48], Proposition 1.2. To this end, note first that the family of functions $\{{P}_{\eta}(x),\eta\in\Gamma,{\min_{i}\eta_{i}>1}\}$, with $P_{\eta}$ as in (4), are a linear basis for the algebra taken as the domain of $L_{\alpha,\theta}$. Cf. [60], Section 2. Recall that these functions are intended as their continuous extension from $\nabla_{\infty}$ to $\overline{\nabla}_{\infty}$, i.e., they are evaluated on $\nabla_{\infty}$ and extended to $\overline{\nabla}_{\infty}$ by continuity. Hence ${P}_{(1)}(x)=1$, which for example implies that $1={P}_{(1)}(x){P}_{(1)}(x)={P}_{(2)}(x)+{P}_{(1,1)}(x)$, from which one finds that ${P}_{(1,1)}(x):=1-{P}_{(2)}(x)$. More generally, we would like to compute $L_{\alpha,\theta}$ on all $P_{\eta}$, which the following lemma allows. Denote by $e_{i}$ the canonical vector in the $i$-th direction, and let $\eta-e_{i}$, when $\eta_{i}=1$, indicate $(\eta_{1},\ldots,\eta_{i-1},\eta_{i+1},\ldots,\eta_{d})$. In the next two Lemmas, we assume for simplicity of exposition that $\eta$ is unranked, which has the only purpose of avoiding to account for multiplicity constants. The actual value of $P_{\eta}$ is unaffected since it is symmetric in the $\eta$ components.
@@
Note that ${P}_{\eta-e_{i}}$ and ${P}_{\eta-e_{i}+e_{j}}$ in the above Lemma are well defined since ${P}_{\eta}$ is symmetric in $\eta$. The next Lemma, key in identifying the dual process, makes use of (9) to show how the operator $L_{\alpha,\theta}$ acts on all functions ${P}_{\eta}$.
@@
$L_{\alpha,\theta}$ is well defined on all ${P}_{\eta}$ in the system of equations (9), recursive on $a_{1}(\eta)$. In particular $L_{\alpha,\theta}1=0$, and if $n=|\eta|$ and $d=l(\eta)$,
@@
This was first proved by [60], Proposition 3.1. In the appendix we provide an independent proof based on (9). Note in particular that $L_{\alpha,\theta}P_{\eta}=0$ when $\eta=(1)$, which uses the extension by continuity from $\nabla_{\infty}$ to $\overline{\nabla}_{\infty}$ recalled in the introduction. ∎
@@
Let $X$ be the diffusion process corresponding to $L_{\alpha,\theta}$, and let $\{\mathcal{D}_{t}\}_{t\geq 0}$ be a continuous-time death process on $\Gamma$, with transition rates
(13) $\lambda_{|\eta|}p^{\downarrow}(\eta,\omega),\quad\quad p^{\downarrow}(\eta,\omega):=\frac{\eta_{i}{a}_{\eta_{i}}(\eta)}{|\eta|},\quad\quad|\eta|>1,$
when $\omega$ is the descending arrangement of $\eta-e_{i}$, and zero elsewhere. Then, for every $\eta\in\Gamma$ and $x\in\overline{\nabla}_{\infty}$, we have
(14) $\mathbb{E}\big{[}g_{\eta}(X(t))|X_{0}=x]=\mathbb{E}\big{[}g_{\mathcal{D}_{t}}(x)|\mathcal{D}_{0}=\eta].$
@@
The proof essentially follows from Lemma 3.2 together with an argument along the lines of that in Section 2 of [5]. Since $g_{\eta}(x)$ is symmetric with respect to $\eta$, from Lemma 3.2 we have
@@
If we now define $A_{\theta}$ to be the operator on the right hand side of (15) acting on $g_{\eta}(x)$ as a function of $\eta$, it is plain that $A_{\theta}$ defines a pure-death process $\mathcal{D}_{t}$ with rates as in the claim. Note that when $|\eta|=1$ the rate is null, as $g_{\eta}\propto 1$ and $L_{\alpha,\theta}g_{\eta}=0$ by Lemma 3.2. Now the fact that $L_{\alpha,\theta}g_{\eta}(x)=A_{\theta}g_{\eta}(x)$ implies (14) follows from Corollary 4.4.13 in [21] in light of the boundedness of $g_{\eta}(x)$ (cf. also Proposition 1.2 in [48]). ∎
Here $\mathcal{D}_{t}$ is the process describing the group sizes in Kingman’s coalescent with mutation at time $t$. The transition rates only depend on the parameter $\theta$ as in the one-parameter model, while the dependence on the second parameter $\alpha$ is only through (11).
The next proposition identifies the transition function of the dual. To this end, let
@@
Here ${\cal H}(\omega|\eta)$ is the probability of obtaining an unlabelled colour configuration $\omega$ when sampling $|\omega|$ balls without replacement uniformly from an urn containing a colour configuration $\eta$. The $|\omega|$ balls may be sampled all at once or one by one.
@@
(20) $q^{\theta}_{\eta\omega}(t):=\mathbb{P}(\mathcal{D}_{t}=\omega|\mathcal{D}_{0}=\eta)={\cal H}(\omega|\eta)d^{\theta}_{|\eta||\omega|}(t),\quad\quad\omega\subset\eta,\quad\quad|\eta|>1,$
and zero otherwise, with $d^{\theta}_{nm}(t)$ for $n\geq m\geq 1$ as in (17). In particular, $\eta=(1)$ is an absorbing state.
@@
$q_{\eta\omega}^{\theta}(t)=$ $\,\mathbb{P}_{\eta}(\mathcal{D}_{t}=\omega)=\mathbb{P}_{\eta}(\mathcal{D}_{t}=\omega|D_{t}=|\omega|)\mathbb{P}_{\eta}(D_{t}=|\omega|)$
@@
Finally, the fact that $\eta=(1)$ is absorbing follows immediately from Theorem 3.3 through the fact that $L_{\alpha,\theta}{P}_{\eta}=0$ when $\eta=(1)$ in Lemma 3.2. ∎
Note that Proposition 3.4 accommodates the full range of values for $\theta>-\alpha$, where $10\leq\alpha<$ 1. In fact, even when $-1<\theta<0$, the transition probabilities in (20) are positive for any $\eta$ such that $|\eta|>1$, through (13) and specifically the fact that $\lambda_{n}>0$ for all $n>1$; cf. (12). The fact that formally $\lambda_{1}<0$ when $-1<\theta<0$ is immaterial, as the dual process is absorbed in $\eta=(1)$.
@@ 4 Two-parameter conditional partition structures
Let $\ell=|\omega|$, and suppose an urn contains a single white ball and $\ell$ non-white balls, whose colours are denoted by $Y^{\ell}=(Y_{1},\cdots,Y_{\ell})$, with $r\leq\ell$ different colours. Define $\pi:Y^{\ell}\to\omega$ to be the function that maps the sample $Y^{\ell}$ into the partition $\omega=(\omega_{1},\cdots,\omega_{r})\in\Gamma_{\ell}$, determined by the equivalence relation on colours, whereby $Y_{i},Y_{j}$ are in the same group if $Y_{i}=Y_{j}$. For every group $\omega_{j}$, one ball is assigned mass $1-\alpha$ and the remaining $\omega_{j}-1$ balls are assigned mass $1$. The single white ball in the urn is assigned mass $\theta+\alpha r$. Balls are then drawn sequentially as follows. If the white ball is drawn, a ball of a new colour is added to the urn and assigned mass $1-\alpha$, and the mass of the white ball is increased by $\alpha$. If a coloured ball is drawn, it is replaced in the urn together with an additional ball of mass $1$ of the same colour. This urn model is an extension of (2) and [45].
When the total number of balls in the urn is $n$, denote the colours of the $n-\ell$ new balls by $X^{n-\ell}=(X_{1},\ldots,X_{n-\ell})$. Let the colour configuration be denoted by $\pi(X^{n-\ell})$ for the new sample and $\pi(X^{n-\ell},Y^{\ell})$ for the combined sample. We are interested in the distribution of these colour configurations conditional on $Y^{\ell}$. To this end, denote by $\text{Dir}(\beta_{1},\ldots,\beta_{r})$ a Dirichlet distribution with parameters $(\beta_{1},\ldots,\beta_{r})$ on the $(r-1)$-dimensional simplex, and let a direct sum notation $\oplus$ indicate an accumulation of the points of two point processes. Finally, denote by $\mathrm{PD}_{\alpha,\theta|\omega}$ a $\mathrm{PD}_{\alpha,\theta}$ distribution conditional on having observed the partition $\omega$, and by $\mathbb{E}_{\alpha,\theta|\omega}$ the corresponding expectation.
@@
Let $\omega\in\Gamma_{\ell},\eta\in\Gamma_{n}$ and $\gamma\in\Gamma_{n-\ell}$. Then $\mathrm{PD}_{\alpha,\theta|\omega}$ equals in distribution the descending arrangement of
(22) $Z_{\ell,r}\mathrm{Dir}(\omega_{1}-\alpha,\ldots,\omega_{r}-\alpha)\oplus(1-Z_{\ell,r})\mathrm{PD}_{\alpha,\theta+\alpha r},$
where $Z_{\ell,r}$ has Beta distribution $\mathrm{Beta}(\ell-r\alpha,\theta+r\alpha)$, independent of everything else. Furthermore,
(23) $\mathbb{P}(\pi(X^{n-\ell})=\gamma|\pi(Y^{\ell})=\omega)=$ $\,\mathbb{E}_{\alpha,\theta|\omega}\big{[}\vec{P}_{\gamma}\big{]},$
$\mathbb{P}(\pi(X^{n-\ell},Y^{\ell})=\eta|\pi(Y^{\ell})=\omega)=$ $\,{\cal H}(\omega|\eta)\frac{\mathbb{E}_{\alpha,\theta}\big{[}\vec{P}_{\eta}\big{]}}{\mathbb{E}_{\alpha,\theta}\big{[}\vec{P}_{\omega}\big{]}},$
@@
Equations (22) and (23) follow from the exchangeability of draws in the above described urn model, together with Corollary 20 in [62]. Consider now a path from $\omega$ to $\eta$. Let $\omega^{i}=\pi(X^{i},Y^{\ell}),1\leq i\leq n-\ell$, $\ell^{i}=|\omega^{i}|$ and $r^{i}$ be the number of colours in $\omega^{i}$. Set $\omega^{0}=\omega$. When $\pi(X^{n-\ell},Y^{\ell})=\eta$, there is a path $\omega=\omega^{0}\subset\cdots\subset\omega^{n-\ell}=\eta$. Let $\chi_{B}(\omega^{i},\omega^{i+1})=a_{\omega^{i}_{k}}(\omega^{i})$ if $\omega^{i+1}$ is obtained from $\omega^{i}$ by adding $1$ to an existing $k$-th component or $\chi_{B}(\omega^{i},\omega^{i+1})=1$ if a new component is added to $\omega^{i+1}$. Then
$\mathbb{P}(\omega^{i+1}|\omega^{i})=\chi_{B}(\omega^{i},\omega^{i+1})\frac{\mathbb{E}_{\alpha,\theta}\big{[}{P}_{\omega^{i+1}}\big{]}}{\mathbb{E}_{\alpha,\theta}\big{[}{P}_{\omega^{i}}\big{]}}=\begin{cases}\displaystyle{\small a_{\omega_{k}^{i}}(\omega^{i})}\frac{\omega_{k}^{i}-\alpha}{\theta+\ell^{i}},\quad&\text{if a coloured ball is drawn,}\\[8.53581pt] \displaystyle\frac{\theta+\alpha r^{i}}{\theta+\ell^{i}},&\text{if the white ball is drawn.}\end{cases}$
@@
$\mathbb{P}(\eta|\omega)=$ $\,\sum_{\omega=\omega^{0}\subset\cdots\subset\omega^{n-\ell}=\eta}\prod_{i=0}^{n-\ell-1}\chi_{B}(\omega^{i},\omega^{i+1})\frac{\mathbb{E}_{\alpha,\theta}\big{[}{P}_{\omega^{i+1}}\big{]}}{\mathbb{E}_{\alpha,\theta}\big{[}{P}_{\omega^{i}}\big{]}}$
$=$ $\,\left(\sum_{\omega=\omega^{0}\subset\cdots\subset\omega^{n-w}=\eta}\prod_{i=0}^{n-\ell-1}\chi_{B}(\omega^{i},\omega^{i+1})\right)\frac{\mathbb{E}_{\alpha,\theta}\big{[}{P}_{\eta}\big{]}}{\mathbb{E}_{\alpha,\theta}\big{[}{P}_{\omega}\big{]}}$
$=$ $\,\frac{a_{1}(\omega)!\cdots a_{\ell}(\omega)!\dim(\omega,\eta)}{a_{1}(\eta)!\cdots a_{n}(\eta)!}\frac{\mathbb{E}_{\alpha,\theta}\big{[}{P}_{\eta}\big{]}}{\mathbb{E}_{\alpha,\theta}\big{[}{P}_{\omega}\big{]}}$
$=$ $\,\frac{\binom{\ell}{\omega}\dim(\omega,\eta)}{\binom{n}{\eta}}\frac{\binom{n}{\eta}\frac{1}{a_{1}(\eta)!\cdots a_{n}(\eta)!}\mathbb{E}_{\alpha,\theta}\big{[}{P}_{\eta}\big{]}}{\binom{\ell}{\omega}\frac{1}{a_{1}(\omega)!\cdots a_{\ell}(\omega)!}\mathbb{E}_{\alpha,\theta}\big{[}{P}_{\omega}\big{]}}={\cal H}(\omega|\eta)\frac{\mathbb{E}_{\alpha,\theta}\big{[}\vec{P}_{\eta}\big{]}}{\mathbb{E}_{\alpha,\theta}\big{[}\vec{P}_{\omega}\big{]}}$
@@
Note now that when $\gamma$ gives the counts in the new sample and $\eta$ the counts in the combined sample, we have $\mathbb{P}(\pi(X^{n-\ell})=\gamma|Y^{\ell})=\mathbb{P}(\pi(X^{n-\ell},Y^{\ell})=\eta|Y^{\ell})$ and therefore
@@
Let $V$ have distribution $\mathrm{PD}_{\alpha,\theta}$ and $Y^{\ell}$ be a sample from the urn. By Lemma 2.1, $\mathbb{P}(\pi(Y^{\ell})=\omega,V\in dy)=\vec{P}_{\omega}(y)\mathrm{PD}_{\alpha,\theta}(dy).$ Since the marginal distribution are $\mathbb{P}(\pi(Y^{\ell})=\omega)=\mathbb{E}_{\alpha,\theta}\big{[}\vec{P}_{\omega}\big{]}$, by Bayes’ theorem
$\mathbb{P}(V\in dy|\pi(Y^{\ell})=\omega)=\frac{\vec{P}_{\omega}(y)}{\mathbb{E}_{\alpha,\theta}\big{[}\vec{P}_{\omega}\big{]}}\mathrm{PD}_{\alpha,\theta}(dy).$
From (23) in Proposition 4.1, $\mathbb{P}(V\in dy|\pi(Y^{\ell})=\omega)=\mathrm{PD}_{\alpha,\theta|\omega}(dy)$, leading to the result. ∎